OpenStudy (anonymous):

The area of a rectangle is 140. The length is one less than three times the width. Find the length and width.

OpenStudy (anonymous):

Let's say that the width=w and the length=l The problems says that l=3w-1. You know that the area of the rectangle is lw (3w-1)(w)=Area The problem says that the area of the rectangle is 140 (3w-1)(w)=140 Solve.

OpenStudy (anonymous):

So I would go 3w-3(w)=140? Where do I go from there?

OpenStudy (anonymous):

(3w-1)(w)=140, actually. Then you would distribute the w..... (do it and report back)

OpenStudy (anonymous):

I am not quite sure how to distribute that?

OpenStudy (anonymous):

@Grazes

OpenStudy (anonymous):

Can you distribute 3(x+2)?

OpenStudy (anonymous):

3x+6

OpenStudy (anonymous):

@Grazes

OpenStudy (anonymous):

Do you know what x*x=?

OpenStudy (anonymous):

Is it 3w-1w=140?

OpenStudy (anonymous):

So basically, you distribute it the same way: w(3w-1) (3w*w)-(1*w) \[3w^2-w=140\]

OpenStudy (anonymous):

Then, you want to move all the terms to one side so you can factor for the roots\[3w^2-w-140=0\] Then, you have 2 choices for factoring: factoring by grouping and the quadratic equation. Which did you learn in class?

OpenStudy (anonymous):

I'm taking this class online as a refresher course. I don't remember which one we did. @Grazes

OpenStudy (anonymous):

Let's use the quadratic equation: